Single payment loan repayment Personal Finance Problem A person bonows S120 to be repaid in 10 years with 11% amaly compo penalty The loan may be d iterest repaid athe end of any earier year wit, no prepayment a. What amount will be due if the loan is repaid at the end of year 22 b. What is the repayment at the end of year 4 a. The amount due i he loan is repaid at the and of yea 2 Round to the mearest en b. The repayment atthe end of year 4 is S□ Runds, the nearest cent ) c. The amount due at the end of the benth year is (Round to the nearest cent )
Answer #1

Principal = $120

Interest Rate = 11%

a.

If repaid at the end of year 2,

Amount = P(1 + r)n

P = Princial

r = interest rate

n = number of years

Amount after year 2 = $147.85

b.

Repayment at the end of year 4,

Amount = P(1 + r)n

Amount after year 4 = $182.17

c.

Repayment at the end of year 10,

Amount = P(1 + r)n

Amount after year 10 = $340.73